# Breaking degeneracies in exoplanetary parameters through self-consistent atmosphere-interior modelling

### v2

@@ Abstract
Context. With a new generation of instruments largely dedicated to exoplanets (i.e. JWST, ELTs, PLATO, Ariel) providing atmospheric spectra and mass/radius measurements for large exoplanet populations, planetary models used to characterise findings are being put to the test.
Aims. We seek to develop a new planetary model, the Heat Atmosphere Density Evolution Solver (HADES), by self-consistently coupling an atmosphere and interior model and comparing its results to currently available findings.
Methods. Atmospheric calculations are conducted under radiative-convective equilibrium, while the interior is based on the most recent and validated ab-initio equations of state. Particular attention is paid to the link by insuring a continuous thermal, gravity and molecular mass profile between the two models.
Results. We apply the model to the database of currently known exoplanets to characterise intrinsic thermal properties. In contrast to previous findings, we show that intrinsic temperatures (Tint) of 200-400 K, increasing with equilibrium temperature, are required to explain the observed radius inflation of hot Jupiters. In addition, we apply our model to perform ”atmosphere-interior” retrievals by Bayesian inference using observed spectra and measured parameters. This allows us to show case the model on the examples of WASP-39 b and 51 Eridani b. For the former we show how the use of spectroscopic measurements can break degeneracies in the atmospheric metallicity (Z) and intrinsic temperature. We derive relatively high values : Z = 14.79Solar and TK, necessary to explain the radius inflation and the chemical composition of WASP-39 b. For the latter we show the importance of using a self-consistent model with the radius being a constrained parameter of the model and of using the age of the host star to break radius and mass degeneracies. Applied to 51 Eridani b we derive a planet mass M MJ and a core mass M ME, suggesting a potential formation by core accretion combined with a ”hot start” scenario.
Conclusions. We conclude that self-consistent atmosphere-interior models efficiently break degeneracies in the structure of both transiting and directly imaged exoplanets. Such tools have a great potential to interpret current and future observations, thereby providing new insights into the formation and evolution of exoplanets.
@@
In recent decades, the discovery of strongly irradiated giant exoplanets (Mayor & Queloz 1995) has opened new avenues of exploration in planetary science. These exoplanets, named hot Jupiters, exhibit strikingly different atmospheres compared to the giant planets in our solar system (e.g., Guillot et al. 1996; Guillot & Showman 2002) rendering them compelling subjects for in-depth investigation. Understanding their unique atmospheric characteristics and their impact on planetary observables has become crucial for advancing our knowledge of exoplanetary systems. Furthermore, the discovery of directly imageable exoplanets (e.g., Chauvin et al. 2004, 2005; Marois et al. 2008, 2010; Lagrange et al. 2009) enables us to explore the early, still-warm stages of giant planets, providing valuable insights into their formation processes.
@@
A better understanding of the gas giants of our solar system has had meaningful consequences when it comes to studying exoplanets. On one hand the difficulty to accurately evaluate the evolution of the ice giants raises questions on internal structures (Scheibe et al. 2019). On another hand, the results from the Juno mission shows the ongoing difficulties in understanding the layout of Jupiter’s planetary core combined with the limitations of current equations of state (e.g., Nettelmann et al. 2021; Idini & Stevenson 2022; Militzer et al. 2022a; Howard et al. 2023). The richness of data from our close neighbors far exceeds that which is currently available from exoplanets, which makes them ideal objects of study to fine tune current modelling tools. However, the growing number of discovered exoplanets represents an ideal test bed for evaluating and exploring the wider uses of these tools and performing statistical analyses.
@@
In addressing these complexities, the integration of comprehensive grids of model atmospheres has played a pivotal role (e.g., Baraffe et al. 2003; Phillips et al. 2020; Marley et al. 2021). Each successive generation of models has aimed to refine and augment various physical aspects, whether through improved interior equations of state (EOS) and disequilibrium chemistry (Phillips et al. 2020) or the exploration of varying C/O ratios (Marley et al. 2021). As emphasized by Marley et al. (2021), the diversity of model grids is essential for their complementary nature and intercomparison, enhancing their overall efficacy. Furthering the development of exoplanet model grids requires the inclusion of stellar irradiation and to improve the coherence between planet interior and atmosphere. This work addresses both issues.
@@
The modelling of exoplanet interiors often faces significant degeneracy, underscoring the importance of integrating interior models with atmospheric properties. This integration is exemplified in Bloot et al. (2023), where atmospheric metallicity is utilized to constrain interior designs. Additionally, Müller & Helled (2021) explore how atmospheric properties contribute to constraining planetary bulk composition. Their study of 51 Eri b showcases the use of observed luminosity, derived metallicity, and synthetic evolution tracks to estimate mass and bulk metallicity. The strength of such an approach lies in its coherent connection of various parameters. In our work, we propose advancing this approach further by integrating a more sophisticated atmospheric model capable of generating synthetic spectra. This advancement grants us the ability to derive atmospheric properties from observations while maintaining constraints imposed by the interior. Guzmán-Mesa et al. (2022) exemplify the strengths of this type of model when extracting the properties of GJ436 b, a sub-Neptune-like planet, through a coupled interior-atmosphere model with spectral retrieval. Sub-Neptunes are inherently degenerate, as highlighted by Valencia et al. (2013). Expanding upon this framework, we extend the approach to encompass Jupiter and super-Jupiter-like planets for both transit and direct imaging studies.
@@
Irradiated planets pose a distinct challenge in predicting their observed radii due to the inflated nature of hot Jupiters (Fortney et al. 2021). These bodies necessitate the consideration of stronger intrinsic thermal fluxes than expected for their system’s age to be accurately modeled. Properly accounting for this phenomenon is crucial not only for 1D model grids but also for 3D general circulation models (GCM). Recent studies have demonstrated the possibility of establishing a connection between the equilibrium temperature of a hot Jupiter and its intrinsic thermal flux, providing a plausible explanation for the observed radius (Thorngren & Fortney 2018; Thorngren et al. 2019). Currently, the two favored hypotheses are ohmic dissipation (Batygin & Stevenson 2010) and advection of heat by the atmospheric circulation (Tremblin et al. 2017). This work aims at addressing this issue by directly coupling atmospheric and interior modellings so as to evaluate the evolution of irradiated planets by resolving the thermal evolution equation across a grid of planets in equilibrium at different intrinsic temperatures.
@@
In the subsequent work, we shall try to add to the models available by producing a self-consistent interior-atmosphere model tying physical parameters together. This will, in turn, allow for observations to better constrain parameters such as mass, intrinsic temperature, metallicity, core mass and radius. To complete this, in section 2 we shall explore the models used and how we link the two into one self-consistent one in order to build grids of models. Literature comparisons will be made to verify the veracity of the results. In section 3, we shall evaluate how exoplanets are affected by stellar irradiation. Section 4, will help us understand the various planetary parameters and how they can lead to degeneracies. We shall explore in section 5 how we can use Bayesian inference tools to infer using the developed model parameters explaining an observed spectrum. Finally in sections 6 and 7, we shall apply the model to infer planetary parameters from observations.
@@ 2.1 Atmosphere model
is a 1D radiative-equilibrium model which calculates fluxes using the two-stream approximation assuming hemispheric closure. Radiative-convective equilibrium is solved assuming that the net flux (radiative + convective) is conserved. The conservation of flux over the grid of pressure levels is solved iteratively using constrained linear inversion method. Thermodynamic quantities in Exo-REM
@@
The vertical chemical profiles for the various species are calculated for a given temperature profile assuming some non-equilibrium chemistry between C-, O- and N- bearing compounds. To do this, we use an analytical formulation based on a comparison of chemical time constants with vertical mixing time from Zahnle & Marley (2014). For this calculation, vertical mixing is parameterised by an eddy mixing coefficient Kzz. To maximise inter-comparability between models, we chose to use a constant value for the Kzz as done by Phillips et al. (2020), we take K (with Kzz in ). Phillips et al. (2020) demonstrate that the eddy mixing coefficient has limited effects on evolution models. We include the option to use equilibrium chemistry. Condensed species that would otherwise have formed clouds are removed.
@@
is parameterized by a gravity at a pressure of 1 bar (g), an intrinsic temperature (Tint), an incoming stellar irradiation (Tirr) and a metallicity ratio (Z) which is expressed as a fraction of solar metallicity using values from Asplund et al. (2009). The intrinsic temperature is derived from an incoming flux at the bottom of the atmosphere model, which serves as a boundary condition, using equation 1.
@@
can evaluate the gravity throughout the atmosphere using the hydrostatic balance equation (HSE). The gravity varies with pressure and altitude. It is hence possible to evaluate how gravity varies at the bottom of different atmospheres as a function of the models parameters and most notably, the stellar irradiation as is shown in Figure 1. Exo-REM
@@ 2.2 Interior model
(Licari 2016; Mazevet et al. 2019, 2022) which solves the hydrostatic equations for a given set of ab initio equations of state. Two regions are considered: an H-He envelope and a core that can be divided into an outer ice and an inner rocky parts. The envelope is considered adiabatic and consists of a mixture of hydrogen and helium at a given mass fraction denoted . The core of mass is considered isothermal. The isentrope is computed by interpolation in an EOS dataset at a constant entropy that is fixed at the outer boundary for a given temperature and pressure. The hydrostatic equations are solved for a number of layers of fixed composition and mass and lead to a radius corresponding the mass of the planet considered. The EOS considered in this work for the envelope are the SCVH (Saumon et al. 1995) to reproduce previous work and the reevaluated CMS ab initio EOS (Mazevet et al. 2022) to produce our results. The water ab initio EOS (Mazevet et al. 2019) is used to describe the core as well as the MgSiO3 EOS equally from Mazevet et al. (2019).
@@
For both H-He EOS, the additive rule for mixing hydrogen and helium is used. Recent work, notably Chabrier & Debras (2021) & Howard & Guillot (2023), have shown that corrections to the ideal mixing should be taken into account. Most notably Chabrier & Debras (2021) suggest that by adapting the hydrogen equation of state, it is possible to take into account the interactions between hydrogen and helium. However, as pointed out in Mazevet et al. (2022), such an ad hoc procedure fold in the intrinsic differences between the ab initio results or the fitting procedure used to produce the EOS. As this introduces an uncontrolled parameter into the evaluation of the H-He EOS, we willingly chose not to adopt this approach.
@@
Research on the internal structure of Jupiter has experienced a surge in recent years, marked by significant contributions from missions such as Juno. By integrating insights from these missions with formation models, there has been a notable increase in the complexity of interior models (Vazan et al. 2016; Venturini et al. 2016; Wahl et al. 2017; Debras & Chabrier 2019; Valletta & Helled 2019; Ormel et al. 2021; Debras et al. 2021; Miguel et al. 2022; Helled et al. 2022; Militzer et al. 2022b; Idini & Stevenson 2022). In order to fit the higher order gravitational moments (above J2), models proposing inhomogeneous interiors with dilute cores and/or super-adiabaticity are being proposed. We however chose to remain with a two layered approach, a homogeneous envelope and a solid core that reduces the number of uncontrolled parameters. This approach can be justified with the latest results of Bloot et al. (2023) In this work, they first show through a comparison between a homogeneous and a non-homogeneous interior model applied to exoplanets that the two models are degenerate with the current observations. They secondly show that the differences in bulk metallicity are not substantial. Thirdly, they found little difference in the estimation of the core mass in most cases. Hence in the rest of this work, we do not consider this effect that is ill constrained with the current observations. It should however be kept in mind that by using a two layers model, we formally only represent one hand member of a set of interior models and neglect that exoplanet cores could, as in the case of Jupiter, be diluted.
@@ 2.3 Joining models
We chose to link in a self-consistent way the atmosphere and the interior. In order to do this, we link on pressure (Plink), temperature (T), average molecular mass () and gravity ().
The linkage pressure (Plink) is a function of the atmosphere’s input properties. We can consider two regimes in the atmosphere, radiative and convective, which can be identified in Figure 1 as the approximately isothermal regions for the former and the adiabatic regions for the latter. Plink corresponds approximately to the radiative-convective boundary. In the cases where there are more than one radiative regions, Plink is defined by the deepest radiative-convective boundary, at higher pressure. To correctly join the interior and the atmosphere model, we need to insure that we join uniquely within the convective region. As the interior model is calculated considering an isentropic profile, as such Plink is in practice either equal to or at a slightly higher pressure than the radiative-convective boundary pressure. Exo-REM
determines convective regions of a given profile by checking that the temperature gradient is greater than the adiabatic gradient. Hence once a convective region is determined, we can use the upper most temperature and pressure values within the given region to run Exoris
. The deeper most convective regions are depicted in Figure 1 by the black boxed zones with crosses marking the linkage pressure. We see as the irradiation temperature decreases below the approximate value of the intrinsic temperature the atmosphere is convective at lower pressures. Once Exoris
has computed a radius for this given pressure temperature point at the chosen planetary mass, the gravity can be evaluated. As discussed in section 2.1, for Exo-REM
, the gravity at 1 bar is an input parameter and the gravity is evaluated using the HSE throughout the atmosphere. Hence the gravity at 1 bar requires adjusting in order to match the gravity at Plink, the top pressure of Exoris
, deeper in the atmosphere. As such, Exo-REM
@@
are run iteratively until the gravity at Plink is matched.
The average molecular mass is linked in a direct way by adapting the mass ratio of helium.
Considering the interior mix as :
Which can be linked to the average molecular mass given by the atmosphere at the Plink using :
The hydrogen fraction is derived from the atmospheric model when considering the mixing ratio of various species. Hence heavier elements found in the atmosphere are also accounted for in the interior envelope. The current formulation paves the way for future improvements by including for example the EOS of water as a proxy for heavier elements.
For a converging model, only a few iterations are required, although this varies across the parameter space. The convergence criterion is set for an error below 0.2 for each parameter.
To complete the analysis presented in this work grids of individual models are constructed along the various physical dimensions (M, Tint, Tirr, Z, Mc). Individual models within the grid have different output values (e.g.,radius, spectrum) along these dimensions but are built without altering these input parameters. Table 2 gives the layout of the grid including the grid limits and the step used.

### v3

# Breaking degeneracies in exoplanetary parameters through self-consistent atmosphere–interior modelling

@@ Abstract
Context. With a new generation of observational instruments largely dedicated to exoplanets (i.e. JWST, ELTs, PLATO, and Ariel) providing atmospheric spectra and mass and radius measurements for large exoplanet populations, the planetary models used to understand the findings are being put to the test.
Aims. We seek to develop a new planetary model, the Heat Atmosphere Density Evolution Solver (HADES), which is the product of self-consistently coupling an atmosphere model and an interior model, and aim to compare its results to currently available findings.
Methods. We conducted atmospheric calculations under radiative-convective equilibrium, while the interior is based on the most recent and validated ab initio equations of state. We pay particular attention to the atmosphere–interior link by ensuring a continuous thermal, gravity, and molecular mass profile between the two models.
Results. We applied the model to the database of currently known exoplanets to characterise intrinsic thermal properties. In contrast to previous findings, we show that intrinsic temperatures (Tint) of 200-400 K —increasing with equilibrium temperature— are required to explain the observed radius inflation of hot Jupiters. In addition, we applied our model to perform ‘atmosphere–interior’ retrievals by Bayesian inference using observed spectra and measured parameters. This allows us to showcase the model using example applications, namely to WASP-39 b and 51 Eridani b. For the former, we show how the use of spectroscopic measurements can break degeneracies in the atmospheric metallicity (Z) and intrinsic temperature. We derive relatively high values of Z = 14.79solar and TK, which are necessary to explain the radius inflation and the chemical composition of WASP-39 b. With this example, we show the importance of using a self-consistent model with the radius being a constrained parameter of the model and of using the age of the host star to break radius and mass degeneracies. When applying our model to 51 Eridani b, we derive a planet mass M MJ and a core mass M ME, suggesting a potential formation by core accretion combined with a ‘hot start’ scenario.
Conclusions. We conclude that self-consistent atmosphere–interior models efficiently break degeneracies in the structure of both transiting and directly imaged exoplanets. Such tools have great potential to interpret current and future observations, thereby providing new insights into the formation and evolution of exoplanets.
@@
In recent decades, the discovery of strongly irradiated giant exoplanets (Mayor & Queloz 1995) has opened new avenues of exploration in planetary science. These exoplanets, named hot Jupiters, exhibit strikingly different atmospheres compared to the giant planets in our Solar System (e.g., Guillot et al. 1996; Guillot & Showman 2002), rendering them compelling subjects for in-depth investigation. Understanding their unique atmospheric characteristics and their impact on planetary observables has become crucial for advancing our knowledge of exoplanetary systems. Furthermore, the discovery of directly imageable exoplanets (e.g. Chauvin et al. 2004, 2005; Marois et al. 2008, 2010; Lagrange et al. 2009) enables us to explore the early, still-warm stages of giant planets, providing valuable insights into their formation processes.
@@
A better understanding of the gas giants of our Solar System has had meaningful consequences when it comes to studying exoplanets. On one hand, the difficulty in accurately evaluating the evolution of the ice giants raises questions regarding their internal structure (Scheibe et al. 2019). On the other, the results from the Juno mission highlight the ongoing difficulty in understanding the layout of Jupiter’s planetary core, combined with the limitations of current equations of state (e.g. Nettelmann et al. 2021; Idini & Stevenson 2022; Militzer et al. 2022a; Howard et al. 2023). We still have far more data for our close neighbours than we do for exoplanets, which makes the former ideal objects of study to fine tune current modelling tools. However, the growing number of discovered exoplanets represents an ideal test bed for evaluating and exploring the wider uses of these tools and for performing statistical analyses.
@@
In addressing these complexities, the integration of comprehensive grids of model atmospheres has played a pivotal role (e.g. Baraffe et al. 2003; Phillips et al. 2020; Marley et al. 2021). Each successive generation of models has been designed to refine and augment various physical aspects, whether through improved interior equations of state (EOS) and disequilibrium chemistry (Phillips et al. 2020) or the exploration of the effects of varying C/O ratios (Marley et al. 2021). As emphasised by Marley et al. (2021), the diversity of model grids is essential for their complementary nature and intercomparison, enhancing their overall efficacy. Furthering the development of exoplanet model grids requires the inclusion of stellar irradiation and improvement of the coherence between planet interior and atmosphere. The present work addresses both issues.
@@
The modelling of exoplanet interiors often faces significant degeneracy, underscoring the importance of integrating interior models with atmospheric properties. This integration is exemplified in Bloot et al. (2023), where atmospheric metallicity is used to constrain interior designs. Additionally, Müller & Helled (2021) explore how atmospheric properties contribute to constraining planetary bulk composition. Their study of 51 Eri b showcases the use of observed luminosity, derived metallicity, and synthetic evolution tracks to estimate mass and bulk metallicity. The strength of such an approach lies in its coherent connection of various parameters. In our work, we propose to advance this approach further by integrating a more sophisticated atmospheric model capable of generating synthetic spectra. This advancement grants us the ability to derive atmospheric properties from observations while maintaining constraints imposed by the interior. Guzmán-Mesa et al. (2022) exemplify the strengths of this type of model when extracting the properties of GJ436 b, a sub-Neptune-like planet, through a coupled interior–atmosphere model with spectral retrieval. Our understand- ing of Sub-Neptunes is plagued by degeneracy, as highlighted by Valencia et al. (2013). Expanding upon this framework, we extend the approach to encompass Jupiter and super-Jupiter-like planets for both transit and direct imaging studies.
@@
When irradiated, predicting the observed radii of hot Jupiters is distinctly challenging due to the inflated nature (Fortney et al. 2021). Accurately modelling these bodies necessitates the consideration of stronger intrinsic thermal fluxes than expected for the ages of the systems they are found in. Properly accounting for this phenomenon is crucial not only for 1D model grids but also for 3D general circulation models (GCMs). Recent studies have demonstrated the possibility of establishing a connection between the equilibrium temperature of a hot Jupiter and its intrinsic thermal flux, providing a plausible explanation for the observed radius (Thorngren & Fortney 2018; Thorngren et al. 2019). Currently, the two favoured hypotheses are ohmic dissipation (Batygin & Stevenson 2010) and advection of heat by atmospheric circulation (Tremblin et al. 2017). With the present work, we aim to address this issue by directly coupling atmospheric and interior modellings so as to evaluate the evolution of irradiated planets by resolving the thermal evolution equation across a grid of planets in equilibrium at different intrinsic temperatures.
@@
In this work, we try to add to the models available by producing a self-consistent interior-atmosphere model that ties physical parameters together. We believe that this will allow better constraint of parameters such as mass, intrinsic temperature, metallicity, core mass, and radius from observations. In section 2, we explore the atmosphere and interior models used and how we link the two into one that is self-consistent in order to build grids of models. We make literature comparisons to verify the veracity of the results. In section 3, we evaluate how exoplanets are affected by stellar irradiation. In Section 4, we discuss the various planetary parameters and how they can lead to degeneracies. In section 5, we explore how Bayesian inference tools can be used with our model to infer parameters that explain an observed spectrum. Finally, in Sections 6 and 7, we show how the model can be applied to infer planetary parameters from observations.
@@ 2.1 Atmosphere model
is a 1D radiative-equilibrium model that calculates fluxes using the two-stream approximation assuming hemispheric closure. Radiative-convective equilibrium is solved assuming that the net flux (radiative + convective) is conserved. The conservation of flux over the grid of pressure levels is solved iteratively using a constrained linear inversion method. Thermodynamic quantities in Exo-REM
@@
The vertical chemical profiles for the various species are calculated for a given temperature profile assuming some non-equilibrium chemistry between C-, O-, and N-bearing compounds. To do this, we use an analytical formulation based on a comparison of chemical time constants with vertical mixing time from Zahnle & Marley (2014). For this calculation, vertical mixing is parameterised by an eddy mixing coefficient, Kzz. To maximise the comparability of the models, we chose to use a constant value for the Kzz, as done by Phillips et al. (2020): we take K (with Kzz in ). Phillips et al. (2020) demonstrate that the eddy mixing coefficient has limited effects on evolution models. We include the option to use equilibrium chemistry. Condensed species that would otherwise have formed clouds are removed.
@@
is parameterised by a gravity at a pressure of 1 bar (g), an intrinsic temperature (Tint), an incoming stellar irradiation (Tirr), and a metallicity ratio (Z) expressed as a fraction of solar metallicity using values from Asplund et al. (2009). The intrinsic temperature is derived from an incoming flux at the bottom of the atmosphere model, which serves as a boundary condition, using equation 1:
@@
can evaluate the gravity throughout the atmosphere using the hydrostatic balance equation (HSE). The gravity varies with pressure and altitude. It is hence possible to evaluate how gravity varies at the bottom of different atmospheres as a function of the model parameters, and most notably the stellar irradiation, as shown in Figure 1. Exo-REM
@@ 2.2 Interior model
(Licari 2016; Mazevet et al. 2019, 2022), which solves the hydrostatic equations for a given set of ab initio equations of state. Two regions are considered: an H-He envelope and a core that can be divided into an outer ice and an inner rocky part. The envelope is considered adiabatic and consists of a mixture of hydrogen and helium at a given mass fraction denoted . The core of mass is considered isothermal. The isentrope is computed by interpolation in an EOS dataset at a constant entropy that is fixed at the outer boundary for a given temperature and pressure. The hydrostatic equations are solved for a number of layers of fixed composition and mass and lead to a radius corresponding to the mass of the planet considered. The EOSs considered in this work for the envelope are the SCVH (Saumon et al. 1995), which is used to reproduce the results of previous work, and the reevaluated CMS ab initio EOS (Mazevet et al. 2022), which is used to produce our results. The water ab initio EOS (Mazevet et al. 2019) is used to describe the core, as is the MgSiO3 EOS from Mazevet et al. (2019).
@@
For both H-He EOSs, the additive rule for mixing hydrogen and helium is used. Recent works, such as those of Chabrier & Debras (2021) and Howard & Guillot (2023), have shown that corrections to the ideal mixing should be taken into account. Most notably, Chabrier & Debras (2021) suggest that by adapting the hydrogen EOS, it is possible to take into account the interactions between hydrogen and helium. However, as pointed out in Mazevet et al. (2022), such an ad hoc procedure incorporates the intrinsic differences arising from either the ab initio results or the fitting procedure used to produce the EOS. As this introduces an uncontrolled parameter into the evaluation of the H-He EOS, we willingly chose not to adopt this approach.
@@
Research on the internal structure of Jupiter has experienced a surge in recent years, marked by significant contributions from missions such as Juno. By integrating insights from these missions with formation models, there has been a notable increase in the complexity of interior models (Vazan et al. 2016; Venturini et al. 2016; Wahl et al. 2017; Debras & Chabrier 2019; Valletta & Helled 2019; Ormel et al. 2021; Debras et al. 2021; Miguel et al. 2022; Helled et al. 2022; Militzer et al. 2022b; Idini & Stevenson 2022). In order to fit the higher-order gravitational moments (above J2), models proposing inhomogeneous interiors with dilute cores and/or superadiabaticity are being proposed. We however chose to remain with a two-layered approach: a homogeneous envelope and a solid core, which reduces the number of uncontrolled parameters. This approach appears to be validated by the latest results of Bloot et al. (2023). Through the comparison of a homogeneous and a non-homogeneous interior model applied to exoplanets, these authors show that the two are degenerate with the current observations. They also show that the differences in bulk metallicity are not substantial, and find little difference between the resultant estimations of the core mass in most cases. Given these findings, and the fact that the homogeneity of the interior is ill constrained with the current observations, we do not consider it in the remainder of this work. It should, however, be kept in mind that by using a two-layer model, we are formally representing only one example within a range of possible interior models and are neglecting the possibility that exoplanet cores could be diluted, as observed in the case of Jupiter.
@@ 2.3 Joining models
We chose to link the atmosphere and the interior in a self-consistent way. In order to do this, we link on pressure (Plink), temperature (T), average molecular mass (), and gravity ().
The linkage pressure (Plink) is a function of the atmosphere’s input properties. We can consider two regimes in the atmosphere, radiative and convective, which can be identified in Figure 1 as the approximately isothermal regions for the former and the adiabatic regions for the latter. Plink approximately corresponds to the radiative–convective boundary. In the cases where there ismore than one radiative region, Plink is defined by the deepest radiative–convective boundary, at higher pressure. To correctly join the interior model to the atmosphere model, we need to ensure that we join uniquely within the convective region. As the interior model is calculated considering an isentropic profile, Plink is in practice either equal to or at a slightly higher pressure than the radiative–convective boundary pressure. Exo-REM
determines the convective regions of a given profile by checking that the temperature gradient is greater than the adiabatic gradient. Once a convective region is determined, we can then use the uppermost temperature and pressure values within the given region to run Exoris
. The deeper, most convective regions are depicted in Figure 1 by the black boxed zones with crosses marking the linkage pressure. As the irradiation temperature decreases below the approximate value of the intrinsic temperature, we see that the atmosphere is convective at lower pressures. Once Exoris
has computed a radius for this given pressure–temperature point at the chosen planetary mass, the gravity can be evaluated. As discussed in section 2.1, for Exo-REM
, the gravity at 1 bar is an input parameter and the gravity is evaluated using the HSE throughout the atmosphere. Therefore, the gravity value at 1 bar needs to be adjusted to ensure it matches the gravity at P, the top pressure point of Exoris, located deeper in the atmosphere. Consequently, Exo-REM
@@
are run iteratively until the gravity at P is matched.
The average molecular mass is linked in a direct way by adapting the mass ratio of helium:
Considering the interior mix as
which can be linked to the average molecular mass given by the atmosphere at the Plink using
The hydrogen fraction is derived from the atmospheric model when considering the mixing ratio of various species. Hence, heavier elements found in the atmosphere are also accounted for in the interior envelope. The current formulation paves the way for future improvements by including, for example, the EOS of water as a proxy for heavier elements.
For a converging model, only a few iterations are required, although this varies across the parameter space. The convergence criterion is set for an error of below 0.2 for each parameter.
To complete the analysis presented in this work, grids of individual models are constructed along the various physical dimensions (M, Tint, Tirr, Z, Mc). Individual models within the grid have different output values (e.g. radius and spectrum) along these dimensions but are built without altering these input parameters. Table 2 gives the layout of the grid, including the grid limits and the step used.